OpenStudy (anonymous):

how many solutions does the system of eqations have. x-y=-4 and y=2 (-4,4) and (0,2)

OpenStudy (anonymous):

A. The system has exactly one solution. B. The system has infinitely many solutions. C. The system has no solution

OpenStudy (anonymous):

(-4,4) and (0,2)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What are these? Solution?

OpenStudy (anonymous):

no that's where the lines are on the graph

OpenStudy (anonymous):

x-y=-4 and y=2 these are the solutions

OpenStudy (anonymous):

?????????????

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

So whatcha got to do is quite simply, notice how one of the equations is simply\[y=2\]well, if you plug that into the other equation\[x-y=-4\]you'll get a resulting x-value, as for how to tell if there is one, none, or infinitely many solutions, first step is to look at the slope of the lines since you're dealing with straight lines. If they have equal slope, they are parallel and will have infinitely many solutions if they are actually the SAME line. They have no solutions if they are parallel with a different b value when you have them in the form\[y=mx+b\]and they have one solution if you get a simple number when you do substitution, which is finding where the lines cross eachother

OpenStudy (anonymous):

If you need a further explanation of anything feel free to ask, i'll try my best

OpenStudy (anonymous):

ok and what would be the steps,

OpenStudy (anonymous):

because I understand what you where saying I just don't get how I would plug it in

OpenStudy (anonymous):

so you have\[y=2\]plug that into your other equation X-Y=-4, so you'd have\[X-Y=-4\]\[X-(2)=-4\]add that 2 to both sides to get the X alone and you have that X=-2, so you know they intersect at one point for sure

OpenStudy (anonymous):

and in fact, if you understand that they are both lines, then that has to be the only intersection

OpenStudy (anonymous):

Unless they have same slope, of course

OpenStudy (anonymous):

okay, so it would be (-2)-y=-4 and (-2)-2=-4 right embryo?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No no no, don't replug in the new X-value that you found, once you find the point of intersection, that is your answer, so your final answer would be something along the lines of, "These lines have one point of intersection at (-2,2), and therefore there is only one solution"

OpenStudy (anonymous):

oh okay so there would be only one solution in the end

OpenStudy (anonymous):

Yes that is correct, it's easy to think about it if you draw a bunch of different lines on a paper, and extend those lines until they cross eachother, if they have the same slope, they can never cross, if they have the same slope and the same y-intercept (so they are the same line) then they have infinitely many solutions, if they have different slopes, they have exactly 1 solution

OpenStudy (anonymous):

okay thatnk you so much for the clear explanation, it really helped and I get it now! :)

OpenStudy (anonymous):

and can you maybe help me with another one?

OpenStudy (anonymous):

and can you maybe help me with another problem?

OpenStudy (anonymous):

Go ahead and make a new post (so close this one and ask as a new question) since i currently have to go somewhere, but if it's still unanswered when i get back (in about 30 minutes) i can take a crack at it

OpenStudy (anonymous):

ok thank you so much!!!! :)